OpenStudy (anonymous):

how do i find unit rate of feet per second, when its 85 miles per 1.5 hrs. Also please explain

OpenStudy (anonymous):

Convert 85 miles to feet (1 miles = 5280 feet) Hours to seconds (1 hours = 3600 seconds)

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

85*5280 = 448800 1.5*3600 = 5400

OpenStudy (anonymous):

what do i do after

OpenStudy (anonymous):

Then divide them out: 448800/5400 = 83.1 feet/sec

OpenStudy (anonymous):

thanks :)